# OpsEval: A Comprehensive IT Operations Benchmark Suite for Large Language Models

### v3

@@ B.4. Performance on Different Quantization Models
Figure 12. LLaMA-2-70B’s performance of different quantization parameters. Both zero-shot and few-shot evaluations have been conducted on Chinese Wired Network Operations test set under the naive setting.
B.5. Performance on Different Languages
Figure 13. LLMs’ few-shot performance on English/Chinese test set (CoT+SC)
@@ B.5. Performance on Different Languages
B.6. Case Study
B.6.1. Case study: Why advanced settings sometimes lack behind
Figure 14. A failed CoT example Generated by InternLM-7B on zero-shot evaluation under the CoT+SC setting in Chinese. Original lines in black are annotated by Chinese version in purple, the dark red and the red text are model’s corresponding output.
Figure 15. A failed 3-shot example Generated by Qwen-7B-Chat on both zero-shot and few-shot evaluations under the naive setting in Chinese.
In certain cases, more advanced evaluation methods surprisingly lead to poorer results. We analyze to understand the potential reasons behind this phenomenon:
  • Some models may respond poorly to the guidance provided by the CoT prompts when required to think step by step, leading to subpar outputs. Figure 14 is one of the examples where CoT failed: the model tested cannot comprehend the idea of thinking step by step. Thus, instead of analyzing each option, it repeated the question and came to its answer directly. Even though the model correctly answered “FTP server” when asked in English, it failed to give the expected option A. This failed case inspires the need for few-shot prompting when applying the CoT method.
  • Few-shot prompts may lead some models to believe that the task involves generating questions rather than answering them, resulting in performance issues. Figure 15 provides an example to the problem mentioned above.
B.6.2. Case study: How Baichuan outperforms in Chinese
@@
Figure 16. A failed English-answering example Generated by Baichuan-13B-Chat on few-shot evaluation under the CoT+SC setting in both English and Chinese.

### v4

@@ B.5 Performance on Different Quantization Models
Figure 12: LLaMA-2-70B’s performance of different quantization parameters. Both zero-shot and few-shot evaluations have been conducted on Wired Network Operations test set under the naive setting.
Figure 12 shows the accuracy of LLaMA-2-70B of different quantization parameters on objective questions, English and Chinese questions respectively. We do both zero-shot and few-shot evaluation with the naive setting.
LLaMA2-70B-Int4 can achieve an accuracy close to LLaMA-2-70B without quantization. Specifically, on English multi-choice questions, the accuracy of the GPTQ model with 4-bit quantization parameters is 3.50% lower in zero-shot evaluation and 0.27% in few-shot evaluation compared to LLaMA-2-70B. As for Chinese questions, the accuracy of LLaMA2-70B-Int4 is 3.67% lower in zero-shot evaluation and 5.18% in few-shot evaluation compared to LLaMA-2-70B. However, LLaMA2-70B-Int3 has a performance degradation that cannot be ignored. On average, the accuracy of LLaMA2-70B-Int3 in English set has a 12.46% degradation compared to LLaMA-2-70B and a 9.30% degradation compared to LLaMA2-70B-Int4.
B.6 Performance on Different Languages
Figure 13: LLMs’ few-shot performance on English/Chinese test set (CoT+SC)
@@ B.6 Performance on Different Languages
B.7 Expert alignment of FAE-Score
Figure 14: Scatter plot and trendline of FAE-Score compared to Expert Evaluation score.
As depicted in Figure 14, the FAE-Score demonstrates a strong positive correlation with Expert Evaluation Score, making it a valuable and effective substitute for automated evaluation.
B.8 Case Study
B.8.1 Failure cases of OpsEval questions
Figure 15: LLMs’ failure cases of OpsEval questions.
Figure 15 illustrates three instances of failure in certain LLMs. The primary reasons for these failures are typically attributed to either a deficiency in domain knowledge or a lack of comprehension of domain knowledge.
B.8.2 Case analysis on QA metrics
Figure 16: Case analysis on QA metrics
Figure 16 illustrates an example where BLEU/ROUGE fail to pick up important keywords and are influenced by the repeated pattern attack, while FAE-Score avoids these two issues.
B.8.3 Case study: Why advanced settings sometimes lack behind
Figure 17: A failed CoT example Generated by InternLM-7B on zero-shot evaluation under the CoT+SC setting in Chinese. Original lines in black are annotated by Chinese version in purple, the dark red and the red text are model’s corresponding output.
Figure 18: A failed 3-shot example Generated by Qwen-7B-Chat on both zero-shot and few-shot evaluations under the naive setting in Chinese.
In certain cases, more advanced evaluation methods surprisingly lead to poorer results. We analyze to understand the potential reasons behind this phenomenon. 1) Some models may respond poorly to the guidance provided by the CoT prompts when required to think step by step, leading to subpar outputs. Figure 17 is one of the examples where CoT failed: the model tested cannot comprehend the idea of thinking step by step. Thus, instead of analyzing each option, it repeated the question and came to its answer directly. Even though the model correctly answered “FTP server” when asked in English, it failed to give the expected option A. This failed case inspires the need for few-shot prompting when applying the CoT method. 2) Few-shot prompts may lead some models to believe that the task involves generating questions rather than answering them, resulting in performance issues. Figure 18 provides an example to the problem mentioned above.
B.8.4 Case study: How Baichuan outperforms in Chinese
@@
Figure 19: A failed English-answering example Generated by Baichuan-13B-Chat on few-shot evaluation under the CoT+SC setting in both English and Chinese.
Appendix C Annotation Guideline for OpsEval Categorization
C.1 Overview
In the OpsEval project, we aim to categorize operational and maintenance tasks within the industry. This categorization process is pivotal for understanding the spectrum of tasks and the required abilities to address them effectively. The process involves two primary steps: automated screening using GPT-4 for initial topic modeling, followed by a manual review process involving domain experts.
C.2 Task Categorization
C.2.1 Objective
To categorize questions into one of eight distinct operational tasks based on industry relevance, task frequency, and significance within operational settings.
C.2.2 Steps
  1. 1.
    Review Initial Categorization: Begin with the insights provided by GPT-4’s topic modeling. Each question has been preliminarily categorized into one or more operational tasks.
  2. 2.
    Understand Task Definitions: Familiarize yourself with the details of the eight distinct tasks outlined in the provided Appendix. Each task has specific criteria and examples to guide your categorization.
  3. 3.
    Assign Tasks: For each question, decide which of the eight tasks it belongs to. A question should be categorized based on its core focus and the operational activity it pertains to.
  4. 4.
    Justification: Briefly justify your choice, especially if a question seems to fit into more than one category. Use the task definitions as a guide to support your decision.
C.2.3 Detailed Task Categorizations
  1. 1.
    General Knowledge: Questions related to foundational concepts and practices in the Ops domain.
  2. 2.
    Fault Analysis and Diagnostics: Questions focusing on identifying and solving discrepancies or faults in systems or networks.
  3. 3.
    Network Configuration: Questions about optimal configurations for network devices to ensure efficient and secure operations.
  4. 4.
    Software Deployment: Questions dealing with the distribution and management of software applications.
  5. 5.
    Monitoring and Alerts: Questions on using monitoring tools to oversee system efficiency and setting up alert mechanisms.
  6. 6.
    Performance Optimization: Questions aimed at enhancing network and system performance.
  7. 7.
    Automation Scripts: Questions involving the creation of scripts to automate processes and reduce manual intervention.
  8. 8.
    Miscellaneous: Questions that do not fit into the above categories or involve elements from multiple categories.
C.2.4 Task Categorization Template
Question ID: Question: [Insert question text here] Assigned Task: Justification: [Provide a brief explanation for the task assignment here]
C.2.5 Example for Task Categorization
Question ID: 001 Question: What steps should be taken to configure a firewall to prevent unauthorized access while allowing legitimate traffic? Assigned Task: Network Configuration Justification: This question specifically asks for optimal configuration strategies for a key network device (firewall) to ensure security and efficient operation, aligning perfectly with the ’Network Configuration’ task.
C.3 Ability Categorization
C.3.1 Objective
To classify questions based on the required cognitive ability to answer them: Knowledge Recall, Analytical Thinking, or Practical Application.
C.3.2 Steps
  1. 1.
    Review Definitions: Read the descriptions of the three abilities in the provided Appendix. Each ability category has distinct characteristics and examples.
  2. 2.
    Evaluate Questions: Assess the cognitive demand of each question. Consider what is primarily required to answer the question effectively: recalling information, analyzing data/situations, or applying knowledge in practical scenarios.
  3. 3.
    Assign Ability Level: Determine the most appropriate ability category for each question. Some questions may seem to require multiple abilities; choose the one that is most critical for addressing the core challenge of the question.
  4. 4.
    Justification: Provide a rationale for your categorization, especially for questions that may not clearly fit into a single category. Refer to the ability definitions to support your categorization.
C.3.3 Detailed Ability Categorizations
  1. 1.
    Knowledge Recall: Requires recognizing and recalling core concepts and foundational knowledge.
  2. 2.
    Analytical Thinking: Demands deeper thought to dissect problems, correlate information, and derive conclusions.
  3. 3.
    Practical Application: Involves applying knowledge or analytical insights to provide actionable recommendations.
C.3.4 Ability Categorization Template
Question ID: Question: [Insert question text here] Assigned Ability: Justification: [Provide a brief explanation for the ability level assignment here]
C.3.5 Example for Ability Categorization
Question ID: 002 Question: How would you optimize the performance of a network experiencing frequent bottlenecks? Assigned Ability: Practical Application Justification: The question requires applying knowledge of network systems and performance optimization techniques to propose specific solutions, hence it falls under ’Practical Application’.
C.4 General Guidelines
  • Consistency: Strive for consistency in your categorization decisions. If similar questions are categorized differently, reassess your choices to ensure they align with the task and ability definitions.
  • Collaboration: When in doubt, discuss challenging questions with fellow experts. Collaboration can help clarify ambiguities and refine the categorization process.
  • Documentation: Keep detailed notes on your decisions, especially for questions that required significant deliberation. This documentation will be valuable for future reference and analysis.
By following these guidelines, you will contribute to a comprehensive and nuanced categorization of operational tasks and required abilities. This effort is crucial for enhancing our understanding of the operational landscape and the diverse skills professionals need to navigate it effectively.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 6.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Section 6.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes]
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes]
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] No, because the reported model performance is significant enough that p-values are not needed to verify the results.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] In Appendix B.3, we present the compute resources used in evaluation.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] We remove data which contains personally identifiable information or offensive content manually. See Section 3.2.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Appendix C.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [No] Since this is a community effort, all participants are voluntary and unpaid.